# IndicSuperTokenizer: An Optimized Tokenizer for Indic Multilingual LLMs

### v1

@@ 6 Conclusion
In this work, we revisit tokenization as a central design choice for multilingual LLMs, focusing on Indic languages that expose the limitations of existing subword methods. Our proposed IndicSuperTokenizer combines linguistically grounded pre-tokenization with a two-stage subword–superword learning process, yielding more compact and semantically faithful vocabularies. Experiments across intrinsic metrics, downstream task performance, ablations, and inference latency demonstrate consistent gains in efficiency, morphological alignment, and deployment cost, establishing tokenization as a key lever for building equitable and scalable multilingual models.
Ethics and Reproducibility Statement
@@
This work focuses on the responsible development of multilingual tokenization methods for Indian languages. We did not collect or utilize any sensitive or Personally Identifiable Information (PII). All external datasets, libraries, and tools employed in this work are appropriately acknowledged through citations. Since the study did not involve personal, medical, or otherwise sensitive information, formal IRB approval was not required. Throughout the process, we aimed to minimize biases that could disadvantage low-resource languages. We provide our exhaustive study to advance the development of inclusive and efficient multilingual language models.
@@
Acknowledgments
We sincerely thank the Krutrim leadership for their continued support in conducting this research. We would also like to thank Gautam Rajeev, Guduru Manoj, Shaharukh Khan and Neel Rachamalla for their help in different parts of this work, and the Research team at Krutrim for their valuable suggestions and insightful discussions.
Appendix A Appendix
A.1 Language Details
@@
Table 18: Mapping of ISO codes to corresponding 22 Indic languages.
@@ B.1 Tokenizer implementation
We based our training code for the tokenizer on the open implementation of SuperBPE [^7] https://github.com/PythonNut/superbpe/tree/main using HuggingFace library (jain2022hugging). We also explored merging tokenizers based on the default priority based BPE in SentencePiece [^8] https://github.com/google/sentencepiece. While we explored implementing the multi-word two stage curriculum in the SentencePiece, we found that it was not trivial. On the other hand, HuggingFace showed issues with the merging strategy. We thus relied on different implementations for different approaches.
@@ B.2 Training details
We provide more details about our training setup as discussed in Section 4.2. Each model was trained for 50B tokens under matched hyperparameters (learning rate, batch size, training steps), aligning FLOPs to isolate tokenizer effects. The evaluation was performed using lm-eval-harness (eval-harness) across standard English benchmarks (MMLU, GSM8K, Winogrande, TriviaQA, HellaSwag, ARC, OpenBookQA, CommonsenseQA, DROP) and Indic benchmarks (IndicCOPA, IndicSentiment, IndicXParaphrase, IndicXNLI (doddapaneni2023leavingindiclanguagebehind), ARC Challenge Indic (sarvamai_arcc_indic), and MILU verma2024milu). We report EM for GSM8K and TriviaQA, F1 for DROP, and Accuracy for other benchmarks. Shot settings were fixed per task: $25$-shot for ARC/ARC Challenge Indic, $10$-shot for HellaSwag, $5$-shot for MMLU, GSM8K, and TriviaQA, and zero-shot for the remainder. This setup allows a direct assessment of how tokenizer design influences pretraining efficiency, semantic representation, and generalization across English and Indic tasks.
Table 19: Pretraining configuration for different tokenizers.
@@
IST LLaMA-3.2 1B 53.18 $5\times 10^{-5}$ 68000 4096 192 200008
@@ C.1 Mismatch Between Loss and Task Performance
Although toknizers, incorporating multi-word often show slightly higher loss(liu2025superbpe) during training compared to models using traditional atomic tokenizers like SentencePiece/BPE, this does not necessarily translate to worse downstream performance. We hypothesize that this is due to two complementary factors. First, the introduction of longer or multi-word tokens such as “to the” or “as well as” increases the number of semantically overlapping candidates, making the model’s prediction space less sharply peaked. This means the model may distribute probability across several plausible completions (e.g., “to”, “to the”, “to be”), thereby lowering the maximum assigned probability to the correct token and inflating the cross-entropy loss. In contrast, other BPE tokenizers often yield only one atomic candidate for such function words, allowing sharper predictions with lower loss. Second, IST tokenizes text into fewer, more meaningful units, so when computing the average loss per token, each mistake contributes more heavily to the total. As a result, although the model learns more compact and generalizable representations, its token-level loss appears higher. This creates a divergence between model loss and real-world task accuracy, indicating that traditional loss curves may underrepresent the representational efficiency and practical utility of compositional tokenizers like IST.
@@ C.2 Morphologically grounded token splitting
We investigate the impact of incorporating morphological information into tokenization for Indic languages (brahma2025morphtok). The approach involves pre-processing text with a morphology analyzer to segment words into morphemes prior to training. This experiment focuses on languages in the Devanagari script.
We compare two variants: Tokenizer A, trained on raw text, and Tokenizer B, trained on morphologically segmented text using morphology analyzer (kunchukuttan2020indicnlp). At inference time, Tokenizer B requires the same pre-processing for consistency. Tokenizer B exhibits more semantically coherent superwords, reflecting meaningful morpheme combinations (Figure 4, 4). This promotes better generalization across related forms and reduces the raw token-to-word ratio, as morpheme-based units are more compressible. Sample outputs (Figures 4, 4) illustrate the contrast between surface-level splits and linguistically aligned segmentations.
Despite these gains, we do not adopt this approach in our final tokenizer. The primary limitation is latency, as the pipeline requires both language identification and morphological analysis. Extending robust analyzers across all Indic languages also introduces engineering overhead and brittle dependencies. Nevertheless, morphology-aware tokenization remains a promising direction if fast, reliable analyzers become widely available.
@@ C.3 More on Glitch tokens
For each tokenizer, we vary $K\in\{10,50,100,150,\dots,400\}$ to select the top-$K$ embeddings closest to a reference vector derived from artificially unused tokens in the vocabulary (land2024fishingmagikarpautomaticallydetecting; geiping2024coercing). For the IST tokenizer, we count the number of multi-word tokens within the top-$K$. For the BPE variant, we count tokens with IDs >180,000>180{,}000, which corresponds to the upper 20K of the vocabulary. Both tokenizers share the first 180K IDs; the difference lies in how the final 20K IDs are utilized: IST allocates this space for frequent multi-word tokens, while the BPE tokenizer continues learning subwords. This design choice allows IST to more effectively utilize the tail of the vocabulary for meaningful units, whereas the BPE tokenizer exhibits overfitting in low-frequency subwords. The trend of these counts across different top-$K$ values is visualized in Figure 5. As $K$ increases, the fraction of multi-word tokens in IST remains low but stable, while the BPE variant consistently shows a higher fraction of under-trained subwords, indicating overfitting in the residual vocabulary space.
Appendix D Extended Results
In the main paper, due to space constraints, we limited the number of tokenizers presented. Here, we provide an extended list including all of our baseline tokenizers.
@@ Appendix D Extended Results
IST-BR 1.86 1.76 2.05 1.75 1.62 1.37 2.20 1.86 1.39 1.39 2.19 1.61 2.29 2.30 1.66 1.67 1.69 1.49 2.68 3.61 1.56 2.12 1.88 1.54
IST 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Appendix E Metrics
E.1 Token-to-Word Ratio
E.2 Bytes-per-token
E.3 Normalized Sequence Length
E.4 Reyni’s Efficiency

### v2

@@ 6 Conclusion
In this work, we revisit tokenization as a central design choice for multilingual LLMs and introduce MUTANT, a principled and general recipe for multilingual tokenizer training. We demonstrate the effectiveness of this framework with MUTANT-Indic, which achieves SOTA performance across Indic languages through systematic analyses of vocabulary size, data distribution, language-specific pre-tokenization, and multi-stage learning. Extensive evaluation across intrinsic efficiency metrics, downstream tasks, ablations, and inference latency, in both pretraining-from-scratch and continual pretraining settings, highlights consistent gains in efficiency and deployment cost. We will additionally publicly release our evaluation framework to support reproducibility and future research.
Limitations
While MUTANT-Indic achieves strong intrinsic efficiency gains and practical throughput improvements, our study has several limitations. First, although we evaluate across 22 Indic languages, English, and code, the analysis is restricted to these language families, and extending the framework to other low-resource or non-Indic languages remains future work. Second, downstream evaluations are limited to models up to 1B parameters due to compute constraints; while prior scaling-law results suggest these gains extrapolate to larger models, we do not validate this empirically. Third, although we analyze morphology-aware pre-tokenization, we do not integrate it into the final tokenizer due to its high inference latency, leaving efficient implementations to future work. Finally, while our evaluation framework standardizes intrinsic metrics, it does not capture all linguistic dimensions relevant to downstream tasks.
Ethical Considerations
@@
This work focuses on the responsible development of multilingual tokenization methods for Indian languages. We did not collect or utilize any sensitive or Personally Identifiable Information (PII). AI tools (like ChatGPT) were used to refine text writing and section phrasing in the paper. All external datasets, libraries, and tools employed in this work are appropriately acknowledged through citations. Since the study did not involve personal, medical, or otherwise sensitive information, formal IRB approval was not required. Throughout the process, we aimed to minimize biases that could disadvantage low-resource languages. We provide our exhaustive study to advance the development of inclusive and efficient multilingual language models.
@@
Acknowledgement
We sincerely thank the Krutrim leadership for their continued support in conducting this research. We would also like to thank Aditya Kallappa, Gautam Rajeev, Guduru Manoj, Shaharukh Khan and Neel Rachamalla for their help in different parts of this work, and the Research team at Krutrim for their valuable suggestions and insightful discussions.
Appendix
Appendix A Language Details
We provide more details about the 22 languages and corresponding scripts considered in our work in Table 14. Additionally, ISO code mapping is also provided in Table 15.
@@ Appendix A Language Details
Table 15: Mapping of ISO codes to corresponding 22 Indic languages.
@@ B.1 Tokenizer implementation
Our training code for the tokenizer is based on the open implementation of SuperBPE [^6] https://github.com/PythonNut/superbpe/tree/main using HuggingFace library (Jain, 2022). We first curate a multilingual corpus of $\sim$50G from OLMo (OLMo et al., 2025), Wikipedia [^7] https://en.wikipedia.org/wiki/, books, PDFs, Common Crawl, and the Sangraha dataset (Khan et al., 2024) where we sample our tokenizer training data from varied sources to get enough representation of the different languages individually. English data was mostly sampled from OLMo and code data from Stackv2 (Lozhkov et al., 2024). The common crawl corpus was language-identified using fastText LID, deduplicated using MinHash, filtered for boilerplate and HTML artifacts, normalized using Unicode NFKC and script-validated to reduce cross-script contamination, similar to FineWeb (Penedo et al., 2024). The final MUTANT-Indic tokenizer uses a shared vocabulary of 200200K tokens, distributed across language scripts and is trained on 1010GB of multilingual high quality data.
In our work, we also explored merging tokenizers based on the default priority based BPE in SentencePiece [^8] https://github.com/google/sentencepiece. While we explored implementing the multi-word two stage curriculum in the SentencePiece, we found that it was not trivial. On the other hand, HuggingFace showed issues with the merging strategy. We thus relied on different implementations for different approaches.
@@ B.2 Training details
We provide more details about our training setup as discussed in Section 4.2. Each model was trained for 50B tokens under matched hyperparameters (learning rate, batch size, training steps), aligning FLOPs to isolate tokenizer effects. The evaluation was performed using lm-eval-harness (Gao et al., 2024) across standard English benchmarks (MMLU, GSM8K, Winogrande, TriviaQA, HellaSwag, ARC, OpenBookQA, CommonsenseQA, DROP) and Indic benchmarks (IndicCOPA, IndicSentiment, IndicXParaphrase, IndicXNLI (Doddapaneni et al., 2023), ARC Challenge Indic (Sarvam AI, 2025), and MILU Verma et al. (2024)). We report EM for GSM8K and TriviaQA, F1 for DROP, and Accuracy for other benchmarks. Shot settings were fixed per task: $25$-shot for ARC/ARC Challenge Indic, $10$-shot for HellaSwag, $5$-shot for MMLU, GSM8K, and TriviaQA, and zero-shot for the remainder. This setup allows a direct assessment of how tokenizer design influences pretraining efficiency, semantic representation, and generalization across English and Indic tasks.
@@
MUTANT-Indic LLaMA-3.2 1B 53.18 $5\times 10^{-5}$ 68000 4096 192 200008
Table 16: Pretraining configuration for different tokenizers.
@@ C.1 Mismatch Between Loss and Task Performance
Although toknizers, incorporating multi-word often show slightly higher loss (Liu et al., 2025a) during training compared to models using traditional atomic tokenizers like SentencePiece/BPE, this does not necessarily translate to worse downstream performance. We hypothesize that this is due to two complementary factors. First, the introduction of longer or multi-word tokens such as “to the" or “as well as" increases the number of semantically overlapping candidates, making the model’s prediction space less sharply peaked. This means the model may distribute probability across several plausible completions (e.g., “to", “to the", “to be"), thereby lowering the maximum assigned probability to the correct token and inflating the cross-entropy loss. In contrast, other BPE tokenizers often yield only one atomic candidate for such function words, allowing sharper predictions with lower loss. Second, MUTANT-Indic tokenizes text into fewer, more meaningful units, so when computing the average loss per token, each mistake contributes more heavily to the total. As a result, although the model learns more compact and generalizable representations, its token-level loss appears higher. This creates a divergence between model loss and real-world task accuracy, indicating that traditional loss curves may underrepresent the representational efficiency and practical utility of compositional tokenizers like MUTANT-Indic.
@@ C.2 Morphologically grounded token splitting
We investigate the impact of incorporating morphological information into tokenization for Indic languages (Brahma et al., 2025). The approach involves pre-processing text with a morphology analyzer to segment words into morphemes prior to training. This experiment focuses on languages in the Devanagari script.
We compare two variants: Tokenizer A, trained on raw text, and Tokenizer B, trained on morphologically segmented text using morphology analyzer  (Kunchukuttan, 2020). At inference time, Tokenizer B requires the same pre-processing for consistency. Tokenizer B exhibits more semantically coherent multi-words, reflecting meaningful morpheme combinations (Figure 5, 5). This promotes better generalization across related forms and reduces the raw token-to-word ratio, as morpheme-based units are more compressible. Sample outputs (Figures 5, 5) illustrate the contrast between surface-level splits and linguistically aligned segmentations.
Despite these gains, we do not adopt this approach in our final tokenizer. The primary limitation is latency, as the pipeline requires both language identification and morphological analysis. For completeness, we evaluated a Hindi morphology-aware tokenizer augmented with a morphological analyzer (Kunchukuttan, 2020) combined with language identification (LID) [^9] https://pypi.org/project/langdetect/. We performed inference on approximately 4-5 MB of Hindi text and measured throughput over 10 runs (with 5 warm-up runs), comparing against our MUTANT-Indic tokenizer. Our tokenizer achieved 194K tokens/sec, whereas the morphology-aware tokenizer achieved 90K tokens/sec, representing a 53.28% reduction in throughput. This slowdown arises entirely from the additional LID and morphology-analysis stages, underscoring the efficiency advantages of our approach even when compared to linguistically informed baselines. Extending robust analyzers across all Indic languages also introduces engineering overhead and brittle dependencies. Nevertheless, morphology-aware tokenization remains a promising direction if fast, reliable analyzers become widely available.
@@ C.3 More on Glitch tokens
For each tokenizer, we vary $K\in\{10,50,100,150,\dots,400\}$ to select the top-$K$ embeddings closest to a reference vector derived from artificially unused tokens in the vocabulary (Land and Bartolo, 2024; Geiping et al., 2024). For the MUTANT-Indic tokenizer, we count the number of multi-word tokens within the top-$K$. For the BPE variant, we count tokens with IDs >180,000>180{,}000, which corresponds to the upper 20K of the vocabulary. Both tokenizers share the first 180K IDs; the difference lies in how the final 20K IDs are utilized: MUTANT-Indic allocates this space for frequent multi-word tokens, while the BPE tokenizer continues learning subwords. This design choice allows MUTANT-Indic to more effectively utilize the tail of the vocabulary for meaningful units, whereas the BPE tokenizer exhibits overfitting in low-frequency subwords. The trend of these counts across different top-$K$ values is visualized in Figure 6. As $K$ increases, the fraction of multi-word tokens in MUTANT-Indic remains low but stable, while the BPE variant consistently shows a higher fraction of under-trained subwords, indicating overfitting in the residual vocabulary space.
C.4 More on Latency and Throughput Evaluation
To obtain reliable latency and throughput measurements, we constructed a 200-example multilingual inference set intended to approximate realistic LLM workloads. The set contains diverse sentence-completion style prompts representative of common generation patterns. We include 20 inputs per language across English and nine major Indic languages, ensuring balanced coverage of script diversity, lexical variation, and syntactic complexity. Table 18 presents the token-length distribution of these examples under both the LLaMA-4 tokenizer and our MUTANT-Indic, allowing a controlled comparison of inference efficiency across tokenization schemes.
Table 18: Token-length statistics for the 200-example inference set. We report min, p75, p90, p99, maximum, and average token lengths.
Tokenizer min p75 p90 p99 max avg
Llama-4 288 805 1157 2583 2869 784
MUTANT-Indic 178 440 541 654 676 379
C.5 Details About Baseline Tokenizers
Tokenizer fertility is shaped by multiple factors including training data distribution, vocabulary construction, and underlying algorithmic choices, yet publicly available documentation on these aspects is often limited. Table 19 summarizes the vocabulary sizes, training methodologies, and any disclosed data distributions for all baseline tokenizers considered in our study.
Tokenizer Vocab Size Training Algorithm / Framework Data Distribution
DeepSeek-R1 128K BPE (undisclosed variant) Not publicly disclosed
Gemma-3 262K SentencePiece 140+ languages
GPT-OSS 200K o200k_harmony (TikToken variant) Not publicly disclosed
LLaMA-3.2-1B 128K BPE / SentencePiece-based Not publicly disclosed
LLaMA-4 200K BPE Not fully disclosed
Mistral-Nemo 131K Tekken tokenizer (TikToken-based) 100+ languages; multilingual + code
Qwen-3 151K Byte-level BPE Not publicly disclosed
Sarvam 68K Not publicly disclosed Not publicly disclosed
Sutra 256K SentencePiece (unigram/BPE hybrid) Balanced multilingual; uniform sampling
Table 19: Summary of baseline tokenizers and publicly available training details.
C.6 Bits-per-Character Evaluation on Multiple Choice Benchmarks
In addition to task-level accuracy, we evaluate the language modeling quality of the models using bits-per-character (BPC) measured on the gold answer continuations of benchmark examples. Following Heineman et al. (2025), we compute BPC as:
$\text{BPC}(y|x)=-\frac{1}{|y|\log 2}\sum_{t=1}^{|y|}\log P(y_{t}\mid x,y_{<t}),$ (1)
where $x$ denotes the prompt, $y$ is the gold continuation (correct answer text), and $|y|$ is the number of character in the continuation. Lower values indicate better modeling of the target continuation.
This metric provides a continuous measure of modeling quality that is independent of the discrete multiple-choice selection. It allows comparison of tokenizers through the likelihood assigned to the correct answers in benchmark datasets, where accuracy is near chance.
English Benchmarks
Table 20 reports BPC values on several English multiple-choice benchmarks. Across these tasks, the model trained with the MUTANT-Indic tokenizer consistently achieves lower BPC compared to the model trained with the LLaMA-4 tokenizer, indicating improved likelihood modeling of the correct answer continuations.
Tokenizer ARC-C ARC-E OBQA HellaSwag CSQA
MUTANT-Indic 1.0284 1.2368 1.5774 0.9867 2.3445
LLaMA-4 1.2575 1.6037 1.8345 0.9983 2.7398
Table 20: Bits-per-character (lower is better) on English multiple-choice benchmarks. ARC-C: ARC-Challenge, ARC-E: ARC-Easy, OBQA: OpenBookQA, CSQA: CommonsenseQA.
Indic Benchmark
We further evaluate BPC on translated ARC-Challenge benchmarks across ten Indic languages. Results are shown in Table 21. The MUTANT-Indic tokenizer generally produces lower BPC values across the evaluated languages, suggesting improved modeling efficiency for the gold answer continuations.
Tokenizer bn gu hi kn ml mr or pa ta te
MUTANT-Indic 1.2747 1.3480 1.3066 1.2113 1.0298 1.3205 1.1975 1.2054 0.9317 1.1535
LLaMA-4 1.2756 1.4275 1.3254 1.3193 1.2264 1.4368 1.3694 1.3935 1.0508 1.2889
Table 21: Bits-per-character (lower is better) on ARC-Challenge across Indic languages.
C.7 Stage-2 Regex Boundary Constraints
Stage 1 pre-tokenization relies on established regex-based segmentation strategies commonly used in multilingual tokenizers (e.g., GPT-2 and LLaMA-style rules). These patterns separate punctuation, normalize whitespace boundaries, and produce stable token units prior to subword learning.
Stage 2 instead applies additional regex constraints that regulate subword merges during tokenizer training. The objective is to enable learning of frequent multi-word expressions while preventing unstable token formations.
Specifically, the Stage 2 regex enforces: (i) sentence boundary preservation, preventing merges across sentence delimiters; (ii) explicit handling of Indic punctuation such as the Devanagari markers ‘|” and ‘||” as hard boundaries; and (iii) controlled segmentation of numeric sequences.
Appendix D Metrics Definitions
Here, we discuss the different intrinsic metrics used in our evaluation framework.
D.1 Token-to-Word Ratio
D.2 Bytes-per-token
D.3 Normalized Sequence Length
D.4 Reyni’s Efficiency
Appendix E Extended Results
In the main paper, due to space constraints, we limited the number of tokenizers presented. Here, we provide an extended list including all of our baseline tokenizers. Table 22 and 23 provides the NSL scores and fertility scores of the other tokenizers considered in our work. Table 25 shows the fertility scores for different normalization techniques in our MUTANT-Indic tokenizer. Additionally, we also compare the performance of two models trained with only Stage-1 and the 2-stage curriculum tokenizer in Table 26.
@@ Appendix E Extended Results
MUTANT-Indic 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
MUTANT-Indic (Stage-1) 1.83 1.74 1.99 1.54 1.56 1.33 2.17 1.83 1.36 1.36 2.15 1.56 2.24 2.27 1.61 1.59 1.65 1.47 2.51 3.60 1.45 2.07 1.83 1.47
MUTANT-Indic (Stage-2) 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 24: Fertility scores ($\downarrow$) comparing our tokenizer after Stage 1 and Stage 2 curriculum. Stage-1 tokenizer corresponds to: cleaned corpus, standard BPE, LLaMA-4 regex, identical vocabulary and data size. The only difference between this baseline and our final MUTANT-Indic is the Stage-2 curriculum phase.
Table 26: Comparison of downstream performance between MUTANT-Indic (Stage-1) and MUTANT-Indic (Stage-2).
English Benchmarks Indic Benchmarks
Dataset MUTANT-Indic-Stage-1 MUTANT-Indic-Stage-2 Dataset MUTANT-Indic-Stage-1 MUTANT-Indic-Stage-1
HellaSwag 0.348 0.357 Indic COPA 0.556 0.556
CommonsenseQA 0.193 0.204 Indic Sentiment 0.557 0.551
OpenBookQA 0.214 0.218 Indic XNLI 0.366 0.346
Winogrande 0.515 0.510 Indic Paraphrase 0.562 0.539
GSM8K 0.021 0.018 MILU (Indic Multi-turn LU) 0.265 0.258
ARC Easy 0.625 0.630 ARC Challenge (Indic) 0.247 0.244
ARC Challenge 0.279 0.292 TriviaQA (Indic) 0.268 0.262
MMLU 0.255 0.249
DROP 0.042 0.036
Average 0.277 0.279 Average 0.403 0.394